# OCCAM: Online Continuous Controller Adaptation with Meta-Learned Models

### v1

@@ Abstract
Control tuning and adaptation present a significant challenge to the usage of robots in diverse environments. It is often nontrivial to find a single set of control parameters by hand that work well across the broad array of environments and conditions that a robot might encounter. Automated adaptation approaches must utilize prior knowledge about the system while adapting to significant domain shifts to find new control parameters quickly. In this work, we present a general framework for online controller adaptation that deals with these challenges. We combine meta-learning with Bayesian recursive estimation to learn prior predictive models of system performance that quickly adapt to online data, even when there is significant domain shift. These predictive models can be used as cost functions within efficient sampling-based optimization routines to find new control parameters online that maximize system performance. Our framework is powerful and flexible enough to adapt controllers for four diverse systems: a simulated race car, a simulated quadrupedal robot, and a simulated and physical quadrotor.
@@ 1 Introduction
Our main contribution is OCCAM, a framework for online controller adaptation to new environments using adaptable learned models of performance metrics. We combine meta-learning with Bayesian regression to train predictive models on simulated data that provide a strong prior on controller performance and are readily fit to data gathered during operation. These models are then used as cost functions to search for optimal control parameters in new environments unseen during training. One of the major advantages is the flexibility of our framework: it can easily be applied to any system with a parameterized controller. To illustrate this, we demonstrate that our method performs online controller adaptation on a simulated race car equipped with a nonlinear control law, a simulated and physical quadrotor micro-aerial-vehicle (MAV) with a model-based geometric controller, and a simulated quadruped with a neural-network locomotion policy.
@@ 2 Related Work
Adaptation in Data-Driven Control: A closely-related line of works combines Reinforcement Learning (RL) for learning low-level control policies with system identification. Short-term histories of states and actions are encoded as latent variables to condition either model-free policies [5, 18, 19, 20, 21] or dynamics models [22], allowing the model to modulate its behavior in different environments. [23, 24, 25, 26] instead learn a repertoire of diverse policies for varying environments. The deep context encoders common in these works might not be accurate when the test system lies outside training distribution. A key idea in our method is to explicitly adapt the model to the observed measurements online.
@@
Adaptive Control: Adaptive control generally seeks to augment the control inputs to a system to cancel out disturbances and align the system’s closed-loop dynamics to those of a desired reference model [35]. Direct approaches such as $\mathcal{L}$1-Adaptive control and Model-reference adaptive control have successfully been applied to robotic aerial vehicles [36, 37, 38, 39], manipulators [40], and bipedal robots [41]. Indirect approaches based on online model learning [2, 3, 42] attempt to estimate the dynamics of the robot in order to compute the control actions. Poor values of the additional tunable adaptive gains in adaptive control can render the controller unstable [43]. It is also unclear how a reference model that is always feasible for the actual system to track might be synthesized, and adaptive control generally requires strong assumptions about model structure and differentiability [38, 39], which our method does not require.
@@ 3 Problem Statement
In this work we address the problem of continuously adapting a controller using a combination of offline simulation data and data gathered from the robot online. The goal of the OCCAM procedure is then to find the control gains at iteration $k$, $g_{k}$, that maximize the reward over some time horizon $\Delta T$:
$g^{*}$ $=\operatorname*{arg\,max}_{g}J(y_{k+\Delta T})=\operatorname*{arg\,max}_{g}J(f_{\theta_{k}}(z_{k:k+\Delta T},u_{k:k+\Delta T},\tau_{k}))$ (1)
Where we let $z_{k}\in\mathbb{R}^{N_{z}}$ denote a vector containing the sensor observations from the system at the $k$’th discrete timestep and $u_{k}\in\mathbb{R}^{N_{u}}$ denote the control inputs applied to the system by the controller. $y_{k+\Delta T}\in\mathbb{R}^{N_{y}}$ denotes a vector of performance measures, such as tracking error, control effort, or speed, computed as a fixed function of the measurements and control history over the horizon $\Delta T$: $y_{k+\Delta T}=f(z_{k:k+\Delta T},u_{k:k+\Delta T},\tau_{k})$. In addition, we specify a reward function, $J$ that maps the vector of performance measures onto a single scalar value: $J(y)\in\mathbb{R}$.
The dynamics of our system are governed by a set of unobserved intrinsic and extrinsic system parameters such as inertial parameters or friction coefficients. We refer to these system parameters collectively as $\theta_{k}$. When generating data in simulation the designer would typically perform domain randomization by sampling values from some subset $\Theta$ to explore the behavior of the system in different regimes.
@@ 4.1 Modeling Controller Performance
The foundation of OCCAM is a predictive model $\hat{f}$ which is designed to predict future performance measures as a function of control parameters. The structure of this prediction model is shown in Figure 2 (left). The recent measurements and control inputs (from a sliding window of size $H$), $z_{k-H:k},u_{k-H:k}$, are fed into a shallow fully connected network that produces a context encoding [5] that captures relevant, observable aspects of the system. This encoded result is combined with an encoding of the current task, $\tau_{k}$ and the proposed control parameters, $g_{k}$ and fed into another fully connected network.
To quickly adapt to new data, our prediction network employs a last linear layer structure [44, 30] wherein the fully connected network produces a matrix $\Phi\in\mathbb{R}^{N_{y}\times N_{b}}$ which is combined with a weight vector $w_{t}\in\mathbb{R}^{N_{b}}$ to produce the final prediction: $\hat{y}_{t}=\Phi w_{t}$. We can think of the columns of the matrix $\Phi=[\phi_{1},...\phi_{N_{b}}]$ as a set of basis functions weighted by $w_{t}$. Here we use $t$ as an index instead of $k$ to denote that the collection of performance measures and adaptation occur at a different rate than the low-level control and sensing. We treat $w_{t}$ as a normally distributed random variable: $w_{t}\sim\mathcal{N}(\mu_{t},\Sigma_{t})$. Because $\hat{y}$ is a linear transformation of $w$, we obtain the following distribution for it: $\hat{y}_{t}\sim\mathcal{N}(\Phi\mu_{t},\;\Phi\Sigma_{t}\Phi^{T})$. We use a Kalman filter with identity dynamics to recursively estimate $w$ as new data, $(\Phi_{t},y_{t})$, is obtained online from the system:
@@
$K$ $=\bar{\Sigma}_{t+1}\Phi\left(\Phi\bar{\Sigma}_{t+1}\Phi^{T}+R\right)^{-1}$
$w_{t+1}$ $=w_{t}+K(y_{t+1}-\Phi w_{t})$
$\Sigma_{t+1}$ $=\left(I-K\Phi\right)\bar{\Sigma}_{t+1}$
Informally adopting standard Kalman Filter terminology, $Q$ is a positive definite matrix that scales with our confidence in keeping $w$ constant, while $R$ captures the error in our model, $\Phi$, and the noise in the measurements of $y$.
@@ 4.2 Meta-Training
             Sample random-size subset $\mathcal{D}_{\text{train}}$ of size $n$ from $\mathcal{D}_{\theta}$;
@@
Our goal is to produce learned basis functions that capture the variation we observe across different system parameters, $\theta$. Note that our approach does not attempt to enumerate or estimate $\theta$ directly, but instead adapts in the latent space of $w$. To achieve this, after pretraining our network for a few epochs using standard gradient descent, we switch to gradient-based meta-learning [31], which consists of a base learner which adapts the base model to data from each “task” (in our case, from systems with varying $\theta$), and an outer loop which adjusts the parameters of the base model based on the base learner’s loss. Our base learner is the recursive update (2) applied to a subset of each system dataset. Because every computation in the Kalman Filter is fully differentiable, once the final weights are computed, the gradient of the loss can be backpropagated through the full computation graph to update the neural network parameters and the Kalman filter parameters $w_{0},\Sigma_{0},Q,R$ using standard automatic differentiation tools such as PyTorch’s autograd engine [45]. Algorithm 1 shows how we use the standard meta-learning loops with our Kalman Filter base learner as the inner loop. Meta-learning $w_{0}$ and $\Sigma_{0}$ is similar in philosophy to Harrison et al. [30] and Bertinetto et al. [46], which also learn the prior distribution for the last-layer weights. However, they do not consider a Kalman Filter formulation for regression, and thus do not include process and measurement covariance matrices $Q$ and $R$.
@@ 4.3 Controller Optimization With Learned Model
We want to use the learned prediction model to solve (1) while also using our knowledge of the uncertainty on the model’s predictions. More specifically, at regular intervals OCCAM seeks to find a choice of control parameters that maximizes the uncertainty-aware reward:
@@
where $\hat{y}_{k+\Delta T}=\Phi(g_{k},\tau,z,u)w_{k}$ is the output of our prediction model, and $\beta$ is a scalar penalty that controls the acceptable confidence bound. While many optimizers could be used, we employ a simple evolutionary modification to random-search in our evaluation settings. In each search iteration, we generate a fixed number of random samples from a uniform distribution $G_{r}=\mathcal{U}(g_{\text{min}},g_{\text{max}})$. We also apply $N_{r}$ random perturbations to the optimal gain vectors found in the $T$ previous trials, generating $T\times N_{r}$ evolutionary search samples $G_{e}$. The full set of samples is thus $G=G_{r}\cup G_{e}$. Each sample is concatenated with the upcoming command, $\tau_{k}$ and system history, $z_{k-H:k},u_{k-H:k}$, and passed through the network to obtain the expected reward. After evaluating all samples, we pick the sample with the highest expected reward.
@@ 5.1 Simulated Evaluation Platforms
To show the flexibility of our framework, we evaluate it on three distinct simulated and one real robotic platform. In the following subsections, we give details on each system and explain how our method is applied. Extensive descriptions of each system and its accompanying controllers are provided in the supplementary material.
2D Race Car with Nonlinear Controller: Our first simulated robotic system is a 2-dimensional car racing around procedurally generated race tracks, modified from OpenAI’s gym environment [47]. The car has unknown mass, engine power, and tire friction as $\theta$, with an encoding of the race track shape as task $\tau$. OCCAM’s goal is to continually optimize the six-dimensional control gains $g$ of a PD steering controller that keeps the car on the centerline of the track, and a proportional acceleration controller ($C$) which contain speed limit parameters that enter nondifferentiably into the controller to set thresholds on the maximum allowable speed of the car in straights and corners. The reward function is a linear combination of the performance measures: inverse average lateral tracking error, inverse average wheelslip, and average speed over the upcoming track segment. OCCAM adapts and selects new gains every time the car traverses $\frac{1}{3}$ of a lap.
@@ 5.2 Model Training and Testing
To generate training data from each simulated platform, we randomly sample $N$ parameter vectors $\theta_{0:N}$. For each $\theta_{i}$, we sample $N_{B}$ random $g$ and $\tau$, drive the system to a random initial dynamical state to collect history $z,u$, and finally roll-out the system with controller $g$ to gather performance metrics $y$. We train a separate basis-function network for each system.
@@ 5.3 Baselines
We compare our method on each test system against three baselines. The first baseline is Reptile [51], a first-order version of the common gradient-based meta-learning method MAML [31]. Reptile uses minibatched gradient descent as the base learner and it only outputs a single point prediction of $y$. For our second baseline, we compare against a GP model whose kernel function is the composition of a network trained only using phase-1 of our method with an RBF kernel, mirroring [17, 29]. We also compare against constant control parameters that were hand-tuned by experts for robust and conservative performance across a variety of commands and system parameters. In our quadrotor example, we additionally consider an $\mathcal{L}$1-Adaptive control [36] baseline adapted from [37], which augments the nominal control with an additional term based on an estimated disturbance vector.
To evaluate the usefulness of adapting the prediction model, we perform an ablation where we do not adapt $w$ from trial to trial and only use $w_{0}$ and $\Sigma_{0}$ in each trial (context-only). In the robotic experiments, the context-only ablation mimics methods that directly estimate latent system parameters from states and actions, such as Kumar et al. [5], since the prediction network only uses the history context encoder and does not utilize our last-layer weight adaptation. To evaluate the usefulness of meta-training, we also run an ablation where last-layer weight adaptation occurs, but the network is not trained using the meta-learning procedure (no-meta).

### v2

@@ Abstract
Control tuning and adaptation present a significant challenge to the usage of robots in diverse environments. It is often nontrivial to find a single set of control parameters by hand that work well across the broad array of environments and conditions that a robot might encounter. Automated adaptation approaches must utilize prior knowledge about the system while adapting to significant domain shifts to find new control parameters quickly. In this work, we present a general framework for online controller adaptation that deals with these challenges. We combine meta-learning with Bayesian recursive estimation to learn prior predictive models of system performance that quickly adapt to online data, even when there is significant domain shift. These predictive models can be used as cost functions within efficient sampling-based optimization routines to find new control parameters online that maximize system performance. Our framework is powerful and flexible enough to adapt controllers for four diverse systems: a simulated race car, a simulated quadrupedal robot, and a simulated and physical quadrotor. The video and code can be found at https://hersh500.github.io/occam.
@@ 1 Introduction
In this paper we present OCCAM, a framework for online controller adaptation to new environments using adaptable learned models of performance metrics. Our main contribution is applying meta-learning and online Bayesian regression to train predictive models on simulated data that provide a strong prior on controller performance and are readily adapted to data gathered during online operation. We show that we can use these models as cost functions to search for optimal control parameters in new environments unseen during training. One of the major advantages is the flexibility of our framework: it can easily be applied to any system with a parameterized controller. We demonstrate that our method performs online controller adaptation on a simulated race car equipped with a nonlinear control law, a simulated and physical quadrotor micro-aerial-vehicle (MAV) with a model-based geometric controller, and a simulated quadruped with a neural-network locomotion policy.
@@ 2 Related Work
Adaptation in Data-Driven Control: A closely-related line of works combines Reinforcement Learning (RL) for learning low-level control policies with system identification. Short-term histories, or contexts, of states and actions are encoded as latent variables to condition either model-free policies [5, 18, 19, 20, 21] or dynamics models [22], allowing the model to modulate its behavior in different environments. [23, 24, 25, 26] instead learn a repertoire of diverse policies for varying environments. The deep context encoders common in these works might not be accurate when the test system lies outside training distribution. A key idea in our method is to explicitly adapt the model to the observed performance online.
@@
Adaptive Control: Adaptive control generally seeks to augment the control inputs to a system to cancel out disturbances and align the system’s closed-loop dynamics to those of a desired reference model [35]. Both indirect and direct adaptive control approaches such as $\mathcal{L}$1-Adaptive control, online model learning, and Model Reference Adaptive Control have successfully been applied to robotic aerial vehicles [36, 37, 38, 39, 40], manipulators [41], and bipedal robots [3, 42]. Poor values of the tunable adaptive gains introduced by adaptive control can render the controller unstable [43]. It also may not be possible for the actual system to track the chosen reference model, and adaptive control generally requires strong assumptions about model structure and differentiability [38, 39], which our method does not require.
@@ 3 Problem Statement
In this work we address the problem of continuously adapting a controller using a combination of offline simulation data and data gathered from the robot online. The goal of the OCCAM procedure is then to find the control gains at iteration $k$, $g_{k}$, that maximize the reward over some upcoming time horizon $\Delta T$ that is chosen by the user:
$g^{*}$ $=\operatorname*{arg\,max}_{g}J(y_{k+\Delta T})=\operatorname*{arg\,max}_{g}J(f_{\theta_{k}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g_{k}},z_{k:k+\Delta T},u_{k:k+\Delta T},\tau_{k}))$ (1)
Where we let $z_{k}\in\mathbb{R}^{N_{z}}$ denote a vector containing the sensor observations from the system at the $k$’th discrete timestep and $u_{k}\in\mathbb{R}^{N_{u}}$ denote the control inputs applied to the system by the controller. $y_{k+\Delta T}\in\mathbb{R}^{N_{y}}$ denotes a vector of performance measures, such as tracking error, control effort, or speed, computed as a fixed function of the control gains, measurements and control history over the horizon $\Delta T$: $y_{k+\Delta T}=f({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g_{k}},z_{k:k+\Delta T},u_{k:k+\Delta T},\tau_{k})$. In addition, we specify a reward function, $J$ that maps the vector of performance measures onto a single scalar value: $J(y)\in\mathbb{R}$.
The dynamics of our system are governed by a set of unobserved intrinsic and extrinsic system parameters such as inertial parameters or friction coefficients. We refer to these system parameters collectively as $\theta_{k}$. When generating data in simulation the designer would typically perform domain randomization by sampling possible values of $\theta$ to explore the behavior of the system in different regimes.
@@ 4.1 Modeling Controller Performance
The foundation of OCCAM is a predictive model $\hat{f}$ which is designed to predict future performance measures as a function of the chosen control parameters. The structure of this prediction model is shown in Figure 2 (left). The recent measurements and control inputs (from a sliding window of size $H$), $z_{k-H:k},u_{k-H:k}$, are fed into a shallow fully connected network that produces a context encoding [5] that captures relevant, observable aspects of the system. This encoded result is combined with an encoding of the current task, $\tau_{k}$ and the proposed control parameters, $g_{k}$ and fed into another fully connected network.
To quickly adapt to new data received from the robot during online deployment, the second fully connected network employs a last linear layer structure [30, 44] wherein the network outputs a matrix $\Phi\in\mathbb{R}^{N_{y}\times N_{b}}$ that is combined with a weight vector $w_{t}\in\mathbb{R}^{N_{b}}$ to produce the final prediction: $\hat{y}_{t}=\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t}}w_{t}$. We can think of the columns of the matrix $\Phi=[\phi_{1},...\phi_{N_{b}}]$ as a set of basis functions weighted by $w_{t}$. Here we use $t$ as an index instead of $k$ to denote that this adaptation occurs at a different rate than the low-level control and sensing, with $\Delta t>\Delta k$. Due to uncertainties in the model and measurements, we model $w_{t}$ as a normally distributed random variable: $w_{t}\sim\mathcal{N}(\mu_{t},\Sigma_{t})$. Because $\hat{y}$ is a linear transformation of $w$, we obtain the following distribution for it: $\hat{y}_{t}\sim\mathcal{N}(\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t}}\mu_{t},\;\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t}}\Sigma_{t}\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t}}^{T})$. We use a Kalman filter with identity dynamics to recursively estimate $w$ as new data, $(\Phi_{t},y_{t})$, is obtained online from the system at test time:
@@
$K$ $=\bar{\Sigma}_{t+1}\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t+1}}\left(\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t+1}}\bar{\Sigma}_{t+1}\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t+1}}^{T}+R\right)^{-1}$
$w_{t+1}$ $=w_{t}+K(y_{t+1}-\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t+1}}w_{t})$
$\Sigma_{t+1}$ $=\left(I-K\Phi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t+1}}\right)\bar{\Sigma}_{t+1}$
Informally adopting standard Kalman Filter terminology, $Q$ is a positive definite matrix that scales with our confidence in keeping $w$ constant, while $R$ captures the error in our model, $\Phi$, and the noise in the measurements of $y$. More information on this formulation can be found in the appendix.
@@ 4.2 Meta-Training
             Sample subset $\mathcal{D}_{\text{train}}$ of random size $n$ from $\mathcal{D}_{\theta}$;
@@
While training our model on offline data, our goal is to produce learned basis functions that capture the variation in performance we observe across different system parameters, $\theta$. Instead of estimating $\theta$ directly, our approach adapts in the latent space of $w$. To achieve this, after pretraining our network for a few epochs using standard gradient descent, we switch to gradient-based meta-learning [31], which consists of a base learner which adapts the base model’s initial parameters to data from different “scenarios”, and an outer loop which adjusts the initial parameters of the base model using the loss achieved by the base learner after adaptation. Our base learner is the recursive Kalman update (2) applied sequentially to $n$ random datapoints gathered from a system with new $\theta$. Because every computation in the Kalman Filter is fully differentiable, once the final weights $w_{n}$ are computed, the gradient of the predictive loss can be backpropagated through the full computation graph to update the neural network parameters and the Kalman filter parameters $w_{0},\Sigma_{0},Q,R$ using standard automatic differentiation tools. Algorithm 1 shows how we use the standard meta-learning loops with our Kalman Filter base learner as the inner loop. Meta-learning $w_{0}$ and $\Sigma_{0}$ is similar in philosophy to Harrison et al. [30] and Bertinetto et al. [45], which also learn the prior distribution for the last-layer weights. However, they do not consider a Kalman Filter formulation for regression, and thus do not include process and measurement covariance matrices $Q$ and $R$.
@@ 4.3 Controller Optimization With Learned Model
When running OCCAM online, we want to use the learned prediction model to solve (1) while also using our knowledge of the uncertainty on the model’s predictions. More specifically, at regular intervals OCCAM seeks to find a choice of control parameters that maximizes the uncertainty-aware reward:
@@
where $\hat{y}_{k+\Delta T}=\Phi(g_{k},\tau,z,u)w_{k}$ is the output of the prediction model, and $\beta$ is a scalar penalty that controls the confidence bound. While many optimizers could be used, we employ a particle filter-based random search in our evaluation settings. In each search iteration, we generate a fixed number of random controller gains from a uniform distribution $G_{r}=\mathcal{U}(g_{\text{min}},g_{\text{max}})$. We also propagate previous samples by applying $N_{r}$ random perturbations to the optimal gains found in the $T$ previous trials, generating $T\times N_{r}$ search samples $G_{e}$. The full set of samples is thus $G=G_{r}\cup G_{e}$. Each sample gain is concatenated with the upcoming command, $\tau_{k}$ and system history, $z_{k-H:k},u_{k-H:k}$, and passed through the network to obtain the expected reward. After evaluating all samples, we pick the sample with the highest expected reward. This optimization approach introduces minimal hyperparameters and is computationally efficient; all search samples can be collated into a single batch and evaluated with a single forward pass through the network.
@@ 5.1 Simulated Evaluation Platforms
To show the flexibility of our framework, we evaluate it on three distinct simulated and one real robotic platform, described below. More details on each system and its accompanying controllers are provided in the supplementary material.
2D Race Car with Nonlinear Controller: Our first simulated robotic system is a 2-dimensional car racing around procedurally generated race tracks, modified from OpenAI’s gym environment [46]. The car has unknown mass, engine power, and tire friction as $\theta$, with an encoding of the race track shape as task $\tau$. OCCAM’s goal is to continually optimize the six-dimensional control gains $g$ of a nonlinear PD steering controller that keeps the car on the centerline of the track, and a proportional acceleration controller that accelerates and brakes the car based on tunable speed thresholds. The speed thresholds improve performance but make the controller nonlinear and nondifferentiable. The reward function is a linear combination of the performance measures: inverse average lateral tracking error, inverse average wheelslip, and average speed over the upcoming track segment. OCCAM adapts and selects new gains every time the car traverses $\frac{1}{3}$ of a lap.
@@ 5.2 Model Training and Testing
To generate training data from each simulated platform, we randomly sample $N$ parameter vectors $\theta_{0:N}$. For each $\theta_{i}$, we sample $N_{B}$ random samples of $g$ and $\tau$, drive the system to a random initial dynamical state to collect history $z,u$, and finally roll-out the system with controller $g$ to gather performance metrics $y$. We train a separate basis-function network for each system.
@@ 5.3 Baselines
We compare our method on each test system against three baselines. The first baseline is Reptile [50], a first-order version of the common gradient-based meta-learning method MAML [31]. Reptile uses minibatched gradient descent as the base learner and it only outputs a single point prediction of $y$. For our second baseline, we compare against a GP model whose kernel function is the composition of our network trained without meta-learning with an RBF kernel, mirroring [17, 29]. We also compare against control parameters that were hand-tuned by experts for robust and conservative performance across a variety of commands and system parameters. In our quadrotor example, we additionally consider an $\mathcal{L}$1-Adaptive control [36] baseline adapted from [37], which augments the nominal control with an additional term based on an estimated disturbance vector.
To evaluate the usefulness of adapting the prediction model, we perform an ablation where we do not adapt $w$ from trial to trial and only use $w_{0}$ and $\Sigma_{0}$ (context-only). This ablation mimics methods that directly estimate latent system parameters from state and action histories, such as Kumar et al. [5], since the prediction network only uses the history context encoder and does not utilize our last-layer weight adaptation. To evaluate the usefulness of meta-training, we also run an ablation where last-layer weight adaptation occurs, but the network is trained without our meta-learning procedure (no-meta).